I need help with this
1/5(2x+5)=x+2/3

Answer by stanbon(75887)

(1/5)(2x+5)=x+(2/3)
------------
(2/5)x + 1 = x + (2/3)
(3/5)x= 1/3
x = (5/3)(1/3)
x = 5/9
============
Cheers,
Stan H.